OpenStudy (anonymous):

Evaluate the discriminant of 25x2 + 350x + 49 = 0. Tell how many solutions the equation has and whether the solutions are real or imaginary. A. one real solution B. two imaginary solutions C. two real solutions D. cannot be determined

OpenStudy (nottim):

wat u think first

OpenStudy (anonymous):

i honestly dont have a clue! ha

OpenStudy (nottim):

oh shoot i forgot too. wats discriminant

OpenStudy (anonymous):

i really dont know

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^{2} - 4ac} }{ 2a }\] the discriminant is \[\sqrt{b^{2} - 4ac}\] because it can be used to discriminate between solutions.

OpenStudy (nottim):

oh. that's relatively simple then. Here's a hint: ax^2-bx-c=0 (minus sign is just there; it doesn't matter). works in conjunction with binarymimic's comment.

OpenStudy (accessdenied):

The discriminant is actually a bit simpler, it's the radicand of that expression: \(b^2 - 4ac\) It is useful because the sign on it tells us whether there is a positive number or negative number under the radical. In the case of a positive, we have two real roots. In the case of a negative, we have two imaginary / complex roots. If we have 0, then we have a single root that is repeated. ((-b +- 0) / (2a))

Parth (parthkohli):

Actually, the discriminant is just \(b^2 - 4ac\).